OpenStudy (anonymous):

6. What is the sum of 3/8 and 1/16? A. 1/6 B. 4/24 C. 1/4 D. 7/16

OpenStudy (anonymous):

D

OpenStudy (anonymous):

make 3/8 have a denominator of 16. So multiply everything by 2 to get 6/16 and add 1/16. Adding fractions just add the top to the top to get 7/16

OpenStudy (anonymous):

3/8+1/16 =(2)3/(2)8+1/16 =6/16+1/16 =7/16 =D.